On January 1,2020, Kendall Inc. began construction of an automated cattle feeder system. The system was finished and ready for use on September 30,2021. Expenditures on the project were as follows:

January 1,2020 200,000

September 1,2020 300,000

December 31,2020 300,000

March 31,2021 300,000

September 30,2021 200,000

Kendall borrowed ₱750,000 on a construction loan at 12% interest on January 1,2020. This loan was outstanding throughout the construction period. The company had ₱4,500,000 in 9% bonds payable outstanding in 2020 and 2021.

Questions:

Based on the above data, answer the following:

1. Average accumulated expenditures for 2020 was

2. Interest capitalized for 2020 was

3. Average accumulated expenditures for 2021 was

Interest capitalized for 2021 was



1
Expert's answer
2021-06-17T13:19:42-0400

Solution:

1.). Average accumulated expenditures for 2020 was ₱300,000

Average accumulated expenditures for 2020 = "(200,000\times \frac{12}{12}) + (300,000\times \frac{4}{12}) + (300,000\times \frac{0}{12})"


= $200,000 + $100,000 + 0

= ₱300,000


2.). Interest capitalized for 2020 was ₱36,000

="(300,000\times 12\%) = 36,000"

Interest capitalized for 2020=  ₱36,000


3.). 1.). Average accumulated expenditures for 2021 was ₱1,036,000

January to September = 800,000 + 36,000 = 836,000

Average accumulated expenditures for 2021 =


"(836,000\times \frac{9}{9}) + (300,000\times \frac{6}{9}) + (300,000\times \frac{0}{9})"


= 836,000 + 200,000 + 0


= ₱1,036,000

 

4.). Interest capitalized for 2021 was ₱86,805


= "(750,000\times 12\%\times \frac{9}{12}) = 67,500"


= "(1,036,000 - 750,000)\times 9\%\times \frac{9}{12}) = 19,305"


Total interest capitalized in 2021 = 67,500 + 19,305 = 86,805


Total interest capitalized in 2021 = ₱86,805